OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)?

OpenStudy (aroub):

7-3/2+2 4/4 1 y=mx+b , m=1, (2,7) 7=1(2)+b 7=2+b 5=b y=x+5

OpenStudy (anonymous):

Not an option :(

OpenStudy (aroub):

ah! give me the options please!

OpenStudy (anonymous):

x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (aroub):

x - y = - 5

OpenStudy (anonymous):

Better :)

OpenStudy (aroub):

hehehe!! :D